OpenStudy (anonymous):

Find the area bounded by the curve and the x-axis for r=theta+sin(2*theta) for 0<=theta<=pi.

OpenStudy (anonymous):

@shamil98

OpenStudy (anonymous):

how comfortable are you with integrals?

OpenStudy (anonymous):

Pretty comfortable, can you just set up the integral for this problem? I'll solve the rest.

OpenStudy (anonymous):

\[\int\limits_{0}^{\pi} x+\sin(2x)dx\]

OpenStudy (anonymous):

do you know what to do from here?

OpenStudy (anonymous):

Isn't the formula A=(1/2)r^2?

OpenStudy (anonymous):

? are you trying to do it without Fundamental theorem of calculus?

OpenStudy (anonymous):

Does it involve the fundamental theorem of calculus?